A suitcase is gently dropped on a conveyor belt moving at 3m/s. If the coefficient of friction between the belt and suitcase is 0.5. Find the displacement of suitcase relative to conveyor belt before the slipping b/w the two is stopped
 

Grade:11

1 Answers

Vikas TU
7 years ago
Friction Force = umg = mg/2
accln. = > g/2 
No vertical accln. as converyor belt is moving horizontally.
From Newton’s second law of eqn.=>
s = 0.5*a*t^2
 and
0 = u – at
t = u/a => 3/0.5g => 6/g
and distance traveled => 0.5 * g/2 * (6/g)^2 => 0.9 meter approx.